OpenStudy (anonymous):

Which fraction is greater: 7/20 or 3/8

OpenStudy (anonymous):

3/8

OpenStudy (kamille):

You your calculator and you will see :) Or if you need to explain you must do steps like this. Have you ever heard about "finding the same denominators'? The smallest same denominator is 40, do you agree? Because 20*2=40 and 5*8=40 too. So, if you want to make it right you need to multiplay 7 and 3 as well, so hows this looks : 7/20=7*2/20*2=14/40 , right? and 3/8=3*5/8*5=15/40 so as you see 14/40 is less than 15/40 And sorry for my English.

OpenStudy (luffingsails):

The key tyler is to find the common denominator as Kamille did. Then you can compare numerators to find out which number is larger.